Write a function that models each variation:
a. x=-1 when y=5, y varies inversely as x.
b.x=3 and y=12 when z=2, z varies directly with y and inversely with x.

Answer by stanbon(75887)

Write a function that models each variation:
a. x=-1 when y=5, y varies inversely as x.
y = k/x
5 = k/-1
k = -5
----
Equation:
y = -5/x
--------------------
b.x=3 and y=12 when z=2, z varies directly with y and inversely with x.
z = k*y/x
-----
2 = k*12/3
2 = 4k
k = 1/2
Equation:
y = (1/2)y/x
-----------------
Cheers,
Stan H.
------------